OpenStudy (minamiraftab):

If sinθ=5/6 and tanθ<0, find cosθ and tanθ

OpenStudy (sshayer):

\[\cos \theta=\sqrt{1-\sin ^2 \theta }\]

OpenStudy (sshayer):

\[\tan \theta<0,\theta~lies~\in~second~quadrant.\]

OpenStudy (sshayer):

\[\cos \theta=-\sqrt{1-\sin ^2 \theta }\]

OpenStudy (minamiraftab):

got it! thank you so much :)

OpenStudy (sshayer):

yw